# Data-Driven Estimation of Quadrotor Motor Efficiency via Residual Minimization

### v1

@@ Abstract
A data-driven framework is proposed for online estimation of quadrotor motor efficiency via residual minimization. The problem is formulated as a constrained nonlinear optimization that minimizes trajectory residuals between measured flight data and predictions generated by a quadrotor dynamics model. A sliding-window strategy enables online estimation, and the optimization is efficiently solved using an iteratively reweighted least squares (IRLS) scheme combined with a primal-dual interior-point method, with inequality constraints enforced through a logarithmic barrier function. Robust z-score weighting is employed to reject outliers, which is particularly effective in motor clipping scenarios where the proposed estimator exhibits smaller spikes than an EKF baseline. Compared to traditional filter-based approaches, the batch-mode formulation offers greater flexibility by selectively incorporating informative data segments. This structure is well-suited for onboard implementation, particularly for applications such as fault detection and isolation (FDI), health monitoring, and predictive maintenance in aerial robotic systems. Simulation results under various degradation scenarios demonstrate the accuracy and robustness of the proposed estimator.
@@ I Introduction
Extensive research on system identification for quadrotors has been conducted to improve control accuracy and state estimation. Most prior works focus on identifying parameters such as mass, inertia, or center of mass, and detecting actuator faults. For example, [1] employed an IMU combined with motor speed commands and an Unscented Kalman Filter (UKF) to estimate mass and inertia using Newton-Euler dynamics, while [2] used a Kalman Filter (KF) combined with Recursive Least Squares (RLS) to estimate the moment of inertia. [3] utilized UKF for estimating inertia, velocity, and angular velocity, and [4] compared least-squares, EKF, and instrumental variable methods for mass estimation. Other approaches include [5], who proposed a robust geometric control framework for quadrotors that incorporates online parameter estimation, and [6] employed UKF and EKF to estimate geometric and inertial parameters. [7] used diffusion models to learn unmodeled effects such as aerodynamic disturbances, mass uncertainty, and sensor noise. For fault detection, methods such as neural networks [8], augmented-state Kalman filters [9], sparse identification and Thau’s observer [10], and nonlinear observers with sliding mode control [11] have been proposed. [12] further demonstrated a data-driven system identification framework using Maximum A Posteriori (MAP) estimation to jointly estimate motor delays and inertial properties. However, relatively little attention has been given to the estimation of motor efficiency for quadrotors, despite its critical impact on energy consumption, stability, and flight time. Among these studies, our previous work [13] introduced a UKF-based estimator integrated into a cooperative multirotor system.
This work proposes a batch-mode, optimization-based framework for estimating quadrotor motor efficiency by minimizing trajectory residuals. While not explored in this study, such formulations also offer potential for integration with learning-based methods. The estimation problem is posed as a least-squares optimization with inequality constraints to ensure the efficiency factors remain within a physically valid range (typically $0\leq\eta\leq 1$). To solve this constrained optimization problem, a primal-dual interior-point method is adopted, augmented with an outer IRLS loop that applies robust z-score weighting to suppress outliers. The inequality constraints are enforced through a logarithmic barrier function, and the interior-point framework was originally introduced by Karmarkar [14] and later extended to nonlinear problems using barrier methods by Fiacco and McCormick [15]. A comprehensive introduction of interior-point techniques, including log-barrier formulations, can be found in Boyd and Vandenberghe [16]. The main contributions of this work are summarized as follows:
@@
    A sliding-window scheme with a primal-dual interior-point method and IRLS reweighting improves efficiency and robustness by rejecting outliers and reducing spikes.
@@
The rest of this paper is organized as follows. Section II describes the quadrotor dynamics and control framework. Section III presents the proposed motor efficiency estimator. Simulation results are given in Section IV, and conclusions are drawn in Section V.
@@ II-B Geometric Tracking Control of Quadrotor
Following the geometric tracking control on $\text{SE}(3)$ developed by Lee et al. [17], a controller is designed to track desired trajectories. The tracking errors are defined as:
@@
where $(\cdot)^{\vee}$ denote the inverse of (5), i.e., $\mathfrak{so}(3)\mapsto\mathbb{R}^{3}$, and $(x_{d},v_{d},R_{d},\Omega_{d})$ are the desired position, velocity, rotation matrix, and angular velocity. Based on (6)–(9), the control inputs, namely the collective thrust $f_{\text{c}}$ and moment $M$, are designed as:
@@
where $k_{x}$, $k_{v}$, $k_{R}$, $k_{\Omega}$, are control gains. The desired attitude $R_{d}$ is constructed by $R_{d}=[b_{2d}\times b_{3d};b_{2d};b_{3d}]\in\text{SO}(3)$, with the direction $b_{3d}$ to be designed as:
$b_{3d}=-\frac{-k_{x}e_{x}-k_{v}e_{v}-mge_{3}+m\ddot{x}_{d}}{\|-k_{x}e_{x}-k_{v}e_{v}-mge_{3}+m\ddot{x}_{d}\|}.$ (12)
Equation (12) defines the direction of the desired total thrust.
@@ II-C Thrust Allocation and Motor Efficiency Modeling
where $d\in\mathbb{R}$ is the distance from the center of mass to each motor, and $c_{\tau f}\in\mathbb{R}$ is the thrust-to-torque coefficient in (15). The motor efficiency matrix $E$ is given by:
@@ II-D Numerical Integration of Quadrotor Dynamics
$\hat{v}_{t+1}(s)$ $=v_{t}+\int_{t}^{t+\Delta t}\bigg[ge_{3}-\frac{f_{\text{c},\text{actual}}Re_{3}}{m}\bigg]d\tau$
$\approx v_{t}+\bigg(ge_{3}-\frac{f_{\text{c},\text{actual}}Re_{3}}{m}\bigg)\Delta t.$ (18)
@@
$\hat{x}_{t+1}(s)$ $=x_{t}+\int_{t}^{t+\Delta t}\bigg[v_{t}+\int_{t}^{\tau}ge_{3}-\frac{f_{\text{c},\text{actual}}Re_{3}}{m}d\tau^{\prime}\bigg]d\tau$
$\approx x_{t}+v_{t}\Delta t+\frac{1}{2}\bigg(ge_{3}-\frac{f_{\text{c},\text{actual}}Re_{3}}{m}\bigg)\Delta t^{2}.$ (19)
@@
$\hat{\Omega}_{t+1}(s)$ $=\Omega_{t}+\int_{t}^{t+\Delta t}\big[J^{-1}(M_{\text{actual}}-\Omega\times J\Omega)\big]d\tau$
$\approx\Omega_{t}+\big[J^{-1}(M_{\text{actual}}-\Omega\times J\Omega)\big]\Delta t.$ (20)
For rotation update, the integration of (3) is given by the matrix exponential:
$\hat{R}_{t+1}(s)=R_{t}\operatorname{exp}([\hat{\Omega}_{t}(s)]_{\times}\Delta t),$ (21)
where (21) can be expanded via a Taylor series:
$\text{exp}([\hat{\Omega}_{t}(s)]_{\times}\Delta t)=\sum_{k=0}^{\infty}\frac{([\hat{\Omega}_{t}(s)]_{\times}\Delta t)^{k}}{k!}.$ (22)
By neglecting higher order terms from (22), (21) can be approximated by:
$\hat{R}_{t+1}(s)\approx R_{t}(I+[\hat{\Omega}_{t}(s)]_{\times}\Delta t).$ (23)
@@ III-A Overview of the Estimation Framework
Fig. 1 illustrates the proposed motor efficiency estimation framework. The estimator operates alongside a geometric tracking controller and takes as input the desired thrust and motor command computed by the controller. Within each sliding window of length $n$, the system collects the measured states including position $x$, velocity $v$, angular velocity $\Omega$, and rotation matrix $R$, together with the control inputs of collective thrust $f_{c}$, and moment $M$. These measurements are compared with model-based predictions, and the resulting residuals are minimized through constrained nonlinear optimization to estimate the efficiency of each motor.
@@ III-B Definition of Trajectory Residuals
Trajectory residuals quantify the discrepancy between measured and predicted states from (18)–(20) and (23) with estimated efficiency vector $s$ in (17). The position, velocity, and angular velocity residuals are defined as:
@@
$\delta\hat{R}(s)$ $=\operatorname{exp}([\hat{\Omega}_{t}(s)]_{\times}\Delta t)$
@@ III-C Formulation of the Optimization Problem
$G=\operatorname{blockdiag}(w_{0}G_{0},\cdots,w_{n}G_{n}),$ (33)
@@ III-D Outlier Rejection and Reweighting
To suppress the influence of abnormal measurements, a median absolute deviation (MAD)–based reweighting scheme is applied at each outer iteration. For the $i$-th segment, the weighted residual energy is
@@
Each residual energy is normalized into a robust z-score:
$z_{i}=\frac{|e_{i}-\tilde{m}|}{\max(\operatorname{\text{MAD}},\epsilon_{\text{min}})}$ (37)
and the corresponding weight is updated by
@@
(35)-(39) jointly apply soft decay and hard rejection, which attenuate moderate outliers while eliminating extreme ones, thereby improving robustness to sensor noise and modeling errors.
@@ III-F Primal-Dual Interior-Point Method
$\min_{s_{t}}\quad F(s_{t})-\frac{1}{\beta}\sum_{i=1}^{8}\text{log}\big(-\phi_{i}(s_{t})\big).$ (46)
@@
$\lambda_{i}\phi_{i}(s_{t})=-\frac{1}{\beta}$ (49)
$\rightarrow$ $-\lambda_{i}\phi_{i}(s_{t})-\frac{1}{\beta}=0.$
@@
Defining primal-dual variable $y=[s,\lambda]^{\top}$ and apply Newton’s method to linearize the KKT residuals:
@@
Solving the Newton system (53)–(54) gives the primal-dual update direction:
@@
where $\kappa\in(0,1)$ and $\epsilon_{\text{tol}}>0$ is a tolerance parameter accounting for noise in the data. The updates used in (57) are:
@@
Equation (58) serves as a convergence indicator and as $\hat{\delta}(s_{t},\lambda)$ decreases towards zero, the iterates approach primal and dual feasibility. The optimization terminates when $\|r_{\text{dual}}\|$ and $\hat{\delta}(s_{t},\lambda)$ falls below specified tolerances. The overall primal-dual interior-point method is summarized in Algorithm 1.
@@ III-G Iteratively Reweighted Least Squares (IRLS)
To enhance robustness, the interior-point method is embedded in an outer IRLS loop. At each iteration, residuals are scored with a MAD-based robust z-score to down-weight outliers and discard extreme segments. The reweighted problem is then solved by the inner primal-dual interior-point method (i.e., Algorithm 1), with convergence achieved in about three outer iterations. The overall procedures of IRLS are presented in Algorithm 2.
@@ 7   Run primal-dual interior-point method for $s^{(k)}_{t}$
Algorithm 2 Iteratively Reweighted Least Squares
@@ III-H Derivatives of Trajectory Residuals
This section derives the partial derivatives of all residuals with respect to the motor efficiency parameters $s$. The derivative of the velocity residual with respect to $s$ is:
@@
The derivative of the position residual with respect to $s$ is:
@@
The derivative of the angular velocity residual with respect to $s$ is:
@@
Based on (62), the derivative of the rotation residual with respect to $s$ is:
@@ III-I EKF Baseline
For comparison, we implement an extended Kalman filter (EKF) as baseline. The EKF state is defined as
@@
This EKF serves as a filter-based baseline for comparison against the proposed batch optimization approach.
@@ IV-B Simulation Results and Transition Robustness
Fig. 2-4 compare the EKF baseline and the proposed estimator under gradual degradation, abrupt fault injection, and combined cases, all with thrust noise level $\sigma_{f}=0.07$. While both methods achieve similar long-term accuracy when calibrated to the same root mean square error (RMSE) and standard deviation levels, their behavior during transitions differs markedly. The EKF exhibits pronounced spikes whenever abrupt changes occur, as incoherent measurements are directly incorporated into the filter updates. In contrast, the proposed IRLS-based estimator effectively down-weights outliers and maintains smooth estimates through transitions, accurately reflecting the true motor efficiency even under sudden faults.
@@ IV-D Estimation and Optimization Convergence
Figure 5: Root mean square error (RMSE) and standard deviation of motor efficiency estimation under various degradation and fault scenarios, comparing the EKF baseline with the proposed method.
@@ V Conclusion
Compared to traditional filter-based methods such as EKF, the proposed batch-mode optimization achieves similar long-term accuracy while exhibiting markedly improved robustness during abrupt transitions. Simulation results under degradation, fault injection, and combined scenarios showed that the estimator avoids the large spikes observed in EKF, particularly in motor clipping cases, by effectively rejecting outliers.

### v2

@@ Abstract
A data-driven framework is proposed for online estimation of quadrotor motor efficiency via residual minimization. The problem is formulated as a constrained nonlinear optimization that minimizes trajectory residuals between measured flight data and predictions generated by a quadrotor dynamics model. A sliding-window strategy enables online estimation, and the optimization is efficiently solved using an iteratively reweighted least squares (IRLS) scheme combined with a primal-dual interior-point method, with inequality constraints enforced through a logarithmic barrier function. Robust z-score weighting is employed to reject outliers, which is particularly effective in motor clipping scenarios where the proposed estimator exhibits smaller spikes than an EKF baseline. Compared to traditional filter-based approaches, the batch-mode formulation allows selective inclusion of data segments via IRLS reweighting and hard-rejection. This structure is well-suited for online estimation and supports applications such as fault detection and isolation (FDI), health monitoring, and predictive maintenance in aerial robotic systems. Simulation results under various degradation scenarios demonstrate the accuracy and robustness of the proposed estimator.
@@ I Introduction
Extensive research on system identification for quadrotors has been conducted to improve control accuracy and state estimation. Most prior works focus on identifying parameters such as mass, inertia, or center of mass, and detecting actuator faults. For example, [15] employed an IMU combined with motor speed commands and an Unscented Kalman Filter (UKF) to estimate mass and inertia using Newton-Euler dynamics, while [11] used a Kalman Filter (KF) combined with Recursive Least Squares (RLS) to estimate the moment of inertia. [1] utilized UKF for estimating inertia, velocity, and angular velocity, and [9] compared least-squares, EKF, and instrumental variable methods for mass estimation. Other approaches include [14], who proposed a robust geometric control framework for quadrotors that incorporates online estimation of inertia parameters, and [16] employed UKF and EKF to estimate geometric and inertial parameters. [5] used diffusion models to learn unmodeled effects such as aerodynamic disturbances, mass uncertainty, and sensor noise. For fault detection, methods such as neural networks [8], augmented-state Kalman filters [17], sparse identification and Thau’s observer [12], and nonlinear observers with sliding mode control [2] have been proposed. [6] further demonstrated a data-driven system identification framework using Maximum A Posteriori (MAP) estimation to jointly estimate motor delays and inertial properties. However, relatively little attention has been given to the estimation of motor efficiency for quadrotors, despite its critical impact on energy consumption, stability, and flight time. Among these studies, our previous work [4] introduced a UKF-based estimator integrated into a cooperative multirotor system.
This work proposes a batch-mode, optimization-based framework for estimating quadrotor motor efficiency by minimizing trajectory residuals. While not explored in this study, such formulations also offer potential for integration with learning-based methods. The estimation problem is posed as a least-squares optimization with inequality constraints to ensure the efficiency factors remain within a physically valid range (typically $0\leq\eta\leq 1$). To solve this constrained optimization problem, a primal-dual interior-point method is adopted, augmented with an outer IRLS loop that applies robust z-score weighting to suppress outliers. The inequality constraints are enforced through a logarithmic barrier function, and the interior-point framework was originally introduced by Karmarkar [10] and later extended to nonlinear problems using barrier methods by Fiacco and McCormick [7]. A comprehensive introduction to interior-point techniques, including log-barrier formulations, can be found in [3] by Boyd and Vandenberghe. The main contributions of this work are summarized as follows:
@@
    A sliding-window scheme with a primal-dual interior-point method and IRLS reweighting improves robustness by rejecting outliers and reducing spikes.
@@
The rest of this paper is organized as follows. Section II describes the quadrotor dynamics and control framework. Section III presents the proposed motor efficiency estimator. Section IV provides the simulation results. Finally, Section V concludes the paper.
@@ II-B Geometric Tracking Control of Quadrotor
Following the geometric tracking control on $\text{SE}(3)$ developed by Lee et al. [13], the controller is designed to track desired trajectories. The tracking errors are defined as:
@@
where $(\cdot)^{\vee}$ denotes the inverse of (5), i.e., $\mathfrak{so}(3)\mapsto\mathbb{R}^{3}$, and $(x_{d},v_{d},R_{d},\Omega_{d})$ are the desired position, velocity, attitude, and angular velocity. Based on (6)–(9), the control inputs of thrust $f_{\text{c}}$ and moment $M$ are designed as:
@@
where $k_{x}$, $k_{v}$, $k_{R}$, $k_{\Omega}$, are control gains.
@@ II-C Thrust Allocation and Motor Efficiency Modeling
where $d\in\mathbb{R}$ is the distance parameter from the center of mass, and $c_{\tau f}\in\mathbb{R}$ is the thrust-to-torque coefficient in (14). The motor efficiency matrix $E$ is given by:
@@ II-D Numerical Integration of Quadrotor Dynamics
$\hat{v}_{t+1}(s)=v_{t}+\bigg(ge_{3}-\frac{f_{\text{c},\text{actual}}Re_{3}}{m}\bigg)\Delta t.$ (17)
@@
$\hat{x}_{t+1}(s)=x_{t}+v_{t}\Delta t+\frac{1}{2}\bigg(ge_{3}-\frac{f_{\text{c},\text{actual}}Re_{3}}{m}\bigg)\Delta t^{2}.$ (18)
@@
$\hat{\Omega}_{t+1}(s)=\Omega_{t}+\big[J^{-1}(M_{\text{actual}}-\Omega\times J\Omega)\big]\Delta t.$ (19)
For the rotation matrix update, a first-order approximation is used:
$\hat{R}_{t+1}(s)=R_{t}(I+[\hat{\Omega}_{t}(s)]_{\times}\Delta t).$ (20)
@@ III-A Overview of the Estimation Framework
Fig. 1 illustrates the proposed motor efficiency estimation framework. The estimator operates alongside the geometric tracking controller and takes as input the desired thrust and motor commands computed by the controller. Within each sliding window of length $n$, the system collects the measured states including position $x$, velocity $v$, angular velocity $\Omega$, and rotation matrix $R$, together with the control inputs of collective thrust $f_{c}$ and moment $M$. These measurements are compared with model-based predictions, and the resulting residuals are minimized through constrained nonlinear optimization to estimate the efficiency of each motor.
@@ III-B Definition of Trajectory Residuals
Trajectory residuals quantify the discrepancy between measured and predicted states from (17)–(20) with estimated efficiency vector $s$ in (16). The position, velocity, and angular velocity residuals are defined as:
@@
$\delta\hat{R}(s)$ $=\exp([\hat{\Omega}_{t}(s)]_{\times}\Delta t)$
@@ III-C Formulation of the Optimization Problem
$G=\operatorname{blockdiag}(w_{0}G_{0},\cdots,w_{n-1}G_{n-1}),$ (30)
@@ III-D Outlier Rejection and Reweighting
To suppress the influence of abnormal measurements, a median absolute deviation (MAD)–based reweighting scheme is applied at each outer iteration. MAD is used as a robust scale estimator because extreme residuals can bias the standard deviation estimate. For the $i$-th segment, the weighted residual energy is
@@
Each residual energy is normalized to a robust z-score:
$z_{i}=\frac{|e_{i}-\tilde{m}|}{\max(\text{MAD},\epsilon_{\text{min}})},$ (34)
where $\epsilon_{\text{min}}>0$ avoids division by zero, and the corresponding weight is updated by
@@
(32)–(36) jointly apply soft decay and hard rejection, which attenuate moderate outliers while eliminating extreme ones, thereby improving robustness to noise and modeling errors.
@@ III-F Primal-Dual Interior-Point Method
$\min_{s_{t}}\quad F(s_{t})-\frac{1}{\beta}\sum_{i=1}^{8}\log\big(-\phi_{i}(s_{t})\big).$ (43)
@@
$-\lambda_{i}\phi_{i}(s_{t})-\frac{1}{\beta}=0.$ (46)
@@
Defining primal-dual variable $y=[s,\lambda]^{\top}$ and applying Newton’s method to linearize the KKT residuals:
@@
Solving the Newton system (51) gives the primal-dual update direction:
@@
where $\kappa\in(0,1)$ and $\epsilon_{\text{tol}}>0$ is a tolerance parameter accounting for data noise. The updates used in (54) are:
@@
Equation (55) indicates convergence: as $\hat{\delta}(s_{t},\lambda)$ decreases towards zero, the iterates approach primal and dual feasibility. The optimization terminates when $\|r_{\text{dual}}\|$ and $\hat{\delta}(s_{t},\lambda)$ falls below specified tolerances. The overall primal-dual interior-point method is summarized in Algorithm 1.
@@ III-G Iteratively Reweighted Least Squares (IRLS)
To enhance robustness, the interior-point method is embedded in an outer IRLS loop. At each iteration, residuals are scored with a MAD-based robust z-score to down-weight outliers and discard extreme segments. The reweighted problem is then solved by the inner primal-dual interior-point method (i.e., Algorithm 1), with convergence achieved in about three outer iterations. The overall sliding-window IRLS estimator for quadrotor motor efficiency is summarized in Algorithm 2.
@@ 7   Run primal-dual interior-point method for $s^{(k)}_{t}$
Algorithm 2 Overall IRLS Estimator
@@ III-H Derivatives of Trajectory Residuals
This section derives the partial derivatives of all residuals with respect to the motor efficiency parameters $s$. The velocity residual derivative with respect to $s$ is:
@@
The position residual derivative with respect to $s$ is:
@@
The angular velocity residual derivative with respect to $s$ is:
@@
Based on (59), the rotation residual derivative with respect to $s$ is:
@@ III-I EKF Baseline
For comparison, we implemented an extended Kalman filter (EKF) as a baseline. The EKF state is defined as
@@
This EKF serves as a filter-based baseline for comparison against the proposed batch optimization approach. No explicit outlier rejection is applied in the EKF baseline.
@@ IV-B Simulation Results and Transition Robustness
Fig. 2-4 compare the EKF baseline and the proposed estimator under gradual degradation, abrupt fault injection, and combined cases, all with thrust noise level $\sigma_{f}=0.07$. While both methods achieve similar accuracy when calibrated to the same root mean square error (RMSE) and standard deviation levels, their behavior during transitions differs markedly. The EKF exhibits pronounced spikes whenever abrupt changes occur, as incoherent measurements are directly incorporated into the filter updates. In contrast, the proposed IRLS-based estimator effectively down-weights outliers and maintains smooth estimates through transitions, accurately reflecting the true motor efficiency even under sudden faults.
@@ IV-D Estimation and Optimization Convergence
Figure 5: Root mean square error (RMSE) and standard deviation of motor efficiency estimation under various degradation and fault scenarios, used to show that the EKF baseline and the proposed method are calibrated to comparable average accuracy.
@@ V Conclusion
Compared to traditional filter-based methods such as EKF, the proposed batch-mode optimization achieves similar accuracy while exhibiting markedly improved robustness during abrupt transitions. Simulation results under degradation, fault injection, and combined scenarios showed that the estimator avoids the large spikes observed in the EKF, particularly in motor clipping cases, by effectively rejecting outliers.